OpenStudy (anonymous):

how to solve; x^2 + 3 = 9x

OpenStudy (anonymous):

set it = to 0

OpenStudy (anonymous):

put it in standard form then factor or use quadratic formula... x^2 + 3 = 9x subtract 9x from bth sides: x^2 - 9x + 3 = 0 this looks like it is not factorable so use the quadratic formula ax^2 + bx + c = 0 can you identify a, b, c in your equation?

OpenStudy (anonymous):

yeah what he said.

OpenStudy (anonymous):

a = 1 b = -9 c = 3

OpenStudy (anonymous):

what now?:)

OpenStudy (anonymous):

ok. with those values, stick it into this formula. the quadratic formula is: \[\large x=\frac {-b \pm \sqrt {b^2-4ac}}{2a}\] you'll get two irrational solutions

OpenStudy (anonymous):

i did all that and now im stuck